OpenStudy (moonlitfate):

(a) Set up a linear system whose solution provides the unknown flow rates. (b) Solve the system for the unknown flow rates

OpenStudy (moonlitfate):

Figure is here.

OpenStudy (michele_laino):

if I apply this rule: "algebraic sum of the flow rates at each node has to be equal to zero" namely the conservation of matter at each node, I can write this algebraic system: \[\Large \left\{ \begin{gathered} 500 + 300 = {x_1} + {x_3} \hfill \\ \hfill \\ {x_1} + {x_4} = 200 + {x_2} \hfill \\ \hfill \\ {x_2} + 100 = 600 + {x_5} \hfill \\ \hfill \\ {x_3} + {x_6} = 400 + 350 \hfill \\ \hfill \\ {x_7} + 600 = {x_4} + {x_6} \hfill \\ \hfill \\ {x_5} + 450 = {x_7} + 400 \hfill \\ \end{gathered} \right.\]